IF CARLA EATS 182 GRAINS OF RICE EVERT DAY, BUT EVERY THIRD DAY SHE EATS TWICE AS MANY. HOW MANY WILL SHE EAT IN 18 DAYS?
Answer by rfer(10432)

18/3=6 days at 364
18-6=12 days at 182
12(182)+6(364)=
2184+2184=4368 grains in 18 days